If a vector A = ax + 2a y + 3az , find its magnitude and direction


1
Expert's answer
2021-09-08T09:47:12-0400

"A=a\hat{x}+2a\hat{y}+3a\hat{z}"


"|A|=\sqrt{a^2+(2a)^2+(3a)^2}=\sqrt{14a^2}\\|A|=\sqrt{14}a"

"\hat{n}=\frac{A}{|A|}"

"\hat{n}=\frac{a\hat{x}+2a\hat{y}+3a\hat{z}}{\sqrt{14}a}"

"\hat{n}=\frac{\hat{x}+2\hat{y}+3\hat{z}}{\sqrt{14}}"